OpenStudy (cutiecomittee123):

Suppose a triangle has the sides a, b, c. Let theta be the side opposite of b. If costheta>0 What must be true a^2+b^2>c^2 a^2+c^2>b^2 a^2+c^2

OpenStudy (owen3):

a^2 + b^2 - 2ab cos theta = c^2 solve for a^2 + b^2 a^2 + b^2 = c^2 + 2ab cos theta now if cos theta >0 , then what can we say , assuming a,b >0

OpenStudy (owen3):

you there?

OpenStudy (cutiecomittee123):

Yes, I am lost. @owen3

OpenStudy (cutiecomittee123):

it wouldn't make sense for any side lengths or angles to be less than zero, so of course they have to be greater than 0.

OpenStudy (michele_laino):

hint: if we apply the theorem of Carnot, to such triangle, we get: \[\Large {b^2} = {a^2} + {c^2} - 2ac\cos \theta \] 

OpenStudy (cutiecomittee123):

Okay I with you, but still lost with the equation part. I am a visual learner so its hard for me to learn when I just see the equations and laws that we need to use to figure this out

OpenStudy (michele_laino):

since \(\cos \theta >0\), then we have: \[\Large 2ac\cos \theta>0 \]

OpenStudy (michele_laino):

so: \[\Large {a^2} + {c^2} - 2ac\cos \theta < {a^2} + {c^2}\]

OpenStudy (cutiecomittee123):